Let X, Y be a independent and identically distributed random variables, each having Poisson distribution with rate parameter λ. Find the probability generating function of W = X + Y and hence the mean and variance of W


1
Expert's answer
2022-05-12T18:22:18-0400

Suppose that random variables "X" and "Y" have Poisson distribution with parameter "\lambda". It means that their densities have the following form: "P(X=k)=\frac{\lambda^ke^{-\lambda}}{k!}," "P(Y=n)=\frac{\lambda^ne^{-\lambda}}{n!}" , where "k,n=0,1,2,...". Compute the density for "W": "P(W=n)=\sum_{l=0}^nP(X=l)P(Y=n-l)=\sum_{l=0}^n\frac{\lambda^{n}e^{-2\lambda}}{l!(n-l)!}=\frac{1}{n!}e^{-2\lambda}\sum_{l=0}^nC_n^l\lambda^n", where "C_n^l=\frac{n!}{l!(n-l)!}". We use the binomial formula and get: "P(W=n)=\frac{2\lambda e^{-2\lambda}}{n!}". Thus, "W" also has a Poisson distribution with parameter: "\mu=2\lambda". It is a known fact. We remind that the probability generating function is defined by: "G(z)=E[z^W]=\sum_{k=0}^{+\infty}P(W=k)z^k=\sum_{k=0}^{+\infty}\frac{\mu^ke^{-\mu}}{k!}z^k=e^{-\mu}\sum_{k=0}^{+\infty}\frac{(\mu z)^k}{k!}=e^{-\mu}e^{\mu z}=e^{\mu(z-1)}"

We used the Taylor series for "e^{\mu z}". We received that the probability generating function has the form: "G(z)=e^{\mu(z-1)}". It is also a known fact. After setting: "z=e^t" we receive: "M(t)=e^{\mu(e^t-1)}". It is a moment-generating function. We remind a known fact that "M^{(n)}(0)=E[W^n]". I.e., "n"-th derivative ("n=0,1,2,..." ) of "M(t)" provides the expectation "E[W^n]." We receive: "M'(t)=\mu e^te^{\mu(e^t-1)}", "M''(t)=\mu e^{t} e^{\mu(e^t-1)}+(\mu e^t)^2e^{\mu(e^t-1)}".

Thus, "E[W]=M'(0)=\mu", "E[W^2]=M''(0)=\mu+\mu^2". The variance is: "Var[W]=E[W^2]-(E[W)]^2=\mu+\mu^2-\mu^2=\mu".

Answer: the probability generating function of "W" is: "G(z)=e^{\mu(z-1)}". The mean is: "E[W]=\mu." The variance is: "Var[W]=\mu."